# Both Text and Images Leaked! A Systematic Analysis of Multimodal LLM Data Contamination

### v1

@@ 4 Experiment
In this section, we demonstrate the practicality of our methodology in verifying contamination of multimodal benchmark datasets across several MLLMs.
4.1 Setup
Models.
We conducted extensive evaluations on eight open-source MLLMs, including LLaVA-1.5-7B Liu et al. (2023a), VILA1.5-3B Lin et al. (2023), Qwen-VL-Chat Bai et al. (2023b), fuyu-8b [^3] https://www.adept.ai/blog/fuyu-8b, idefics2-8b Laurençon et al. (2024), Phi-3-vision-128k-instruct Abdin et al. (2024), Yi-VL-6B AI et al. (2024), InternVL2-8B Chen et al. (2023; 2024b), as well as three proprietary MLLMs: GPT-4o OpenAI (2023), Gemini-1.5-Pro Reid et al. (2024), and Claude-3.5-Sonnet [^4] https://www.anthropic.com/news/claude-3-5-sonnet.
Benchmark Datasets.
Our analysis leverages two multi-choice datasets: ScienceQA Lu et al. (2022) and MMStar Chen et al. (2024a), along with three caption datasets: COCO-Caption2017 Lin et al. (2015), NoCaps Agrawal et al. (2019), and Vintage [^5] https://huggingface.co/datasets/SilentAntagonist/vintage-artworks-60k-captioned. MMStar and Vintage, owing to their recent inception, serve to contrast contamination levels with other datasets. We randomly selected 2000 and 1340 samples from ScienceQA’s training and test sets, respectively, with 1000 samples from the other datasets. Given the unavailability of public test labels for COCO-Caption2017 and NoCaps, we used their validation sets.
4.2 Main Results
@@
Metric CR PCR $\Delta$ IL CR PCR $\Delta$ IL CR PCR $\Delta$ IL
@@
Claude-3.5-Sonnet 70.3 65.0 -5.3 67.3 64.9 -2.4 36.3 36.4 0.1 15.9
Table 1: Comparison of MLLMs’ performance on different multi-choice datasets. Bold values indicate the most significant $\Delta$ or $IL$. “–” denotes that $\Delta$ is significant so that $IL$ will not be calculated.
@@
Metric CR PCR $\Delta$ IL CR PCR $\Delta$ IL CR PCR $\Delta$ IL
@@
idefics2-8b 43.5 42.3 -1.2 42.6 37.5 -5.1 18.5 17.0 -1.5
Phi-3-vision-128k-instruct 38.8 39.3 0.5 19.4 36.9 33.3 -3.6 17.4 11.7 -5.7
Yi-VL-6B 43.9 43.3 -0.6 19.4 37.2 36.1 -1.1 3.3 4.2 0.9 2.8
InternVL2-8B 53.3 51.9 -1.4 48.0 46.2 -1.8 28.0 28.7 0.7 18.8
@@
GPT-4o 58.1 54.4 -3.7 54.2 55.1 0.9 19.4 36.3 38.4 2.1 20.1
Gemini-1.5-Pro 57.5 55.3 -2.2 51.2 52.0 0.8 18.7
Claude-3.5-Sonnet 53.7 51.0 -2.7 50.8 51.5 0.7 20.0 35.2 33.0 -2.2 21.3
Table 2: Comparison of MLLMs’ performance on different caption datasets. We have not detected the contamination of Gemini-1.5-Pro on Vintage yet.
@@
As shown in Table 1, for the training set of ScienceQA, Claude-3.5-Sonnet exhibited a significant $\Delta$ of -5.3, indicating extensive potential contamination. VILA1.5-3B’s high IL suggests that the model may have encountered some benchmark instances during training. For the test set of ScienceQA, all models except Claude-3.5-Sonnet showed insignificant $\Delta$, indicating a low likelihood of large-scale leakage from the benchmark. However, the significant IL from fuyu-8b may indicate that the model has seen some benchmark instances or similar training data.
It is noteworthy that even the latest benchmarks may still be contaminated. For MMStar, fuyu-8b showed a significant $\Delta$, and Claude-3.5-Sonnet exhibited a high IL. Since the data sources for MMStar are older datasets, which have been filtered for image-text relevance, the experimental results align with expectations.
@@
From Table 2, we observe that for the validation sets of COCO-Caption2017 and NoCaps, more than half of the models exhibited significant performance drops due to perturbations, indicating that these two benchmarks may face serious leakage issues. For the training set of Vintage, idefics2-8b showed a significant $\Delta$, and Claude-3.5-Sonnet had the highest IL.
@@
5 Intentional Contamination
This section utilizes intentional contamination to address three research questions:
RQ1: Can MM-Detect effectively identify contamination?
RQ2: What is MM-Detect’s sensitivity level?
RQ3: Does training set leakage result in evaluation bias?
To tackle these questions, we followed the recipe of LLaVA and train multiple 7B parameter models on intentionally contaminated training data from downstream task examples during the visual instruction tuning phase, subsequently evaluating their contamination degree.
5.1 MM-Detect is An Effective Detector
We reproduced the LLaVA-1.5-7B model experiment to obtain a control model, LLaVA-1.5-7B-no-cont, which is devoid of intentional data contamination. Subsequently, we randomly incorporated 2000 training samples from ScienceQA and 1000 validation samples from NoCaps into $D_{\text{tuning}}$, where $D_{\text{tuning}}$ denotes the data used by LLaVA for visual instruction tuning. This data was used to train a contaminated model, LLaVA-1.5-7B-cont, for comparison with LLaVA-1.5-7B-no-cont.
Models ScienceQA Train Set NoCaps Val. Set
CR PCR $\Delta$ CR PCR $\Delta$
LLaVA-1.5-7B-cont 72.9 67.9 -5.0 38.2 32.8 -5.4
LLaVA-1.5-7B-no-cont 61.8 61.2 -0.6 33.0 32.1 -0.9
Table 3: Detection results after actively contaminating the model with the ScienceQA training set and NoCaps validation set, showcasing the effectiveness of our method in accurately identifying contamination.
As depicted in Table 3, the inclusion of contaminated data during the training phase increases the model’s familiarity with the data and its sensitivity to perturbations. This leads to an average CR increase of 8.2%, an average PCR increase of 3.7%, and an average $\Delta$ decrease of 4.5%. These changes illustrate our successful detection of data contamination.
5.2 MM-Detect is Sensitive to Contamination Degrees
Figure 4: MM-Detect captures the increasing contamination levels of models on ScienceQA (test set) and reflects them in the atomic metrics.
To investigate the sensitivity of MM-Detect, We selected 1340 examples from the ScienceQA test set to train three models: a fully contaminated model (LLaVA-1.5-7B-cont-100%) incorporating all 1340 examples into $D_{\text{tuning}}$, a partially contaminated model (LLaVA-1.5-7B-cont-50%) with 670 examples mixed in, and a minimally contaminated model (LLaVA-1.5-7B-cont-10%) with 134 examples mixed in. This approach demonstrates our method’s robustness in reflecting various degrees of model contamination.
Figure 4 reveals our testing’s high sensitivity. As contamination degrees rise from 10% to 50%, there’s a significant increase of 8.7% in CR, 7.3% in PCR, and a decrease of 1.4% in $\Delta$. When contamination escalates from 50% to 100%, CR again significantly rises by 5.3%, PCR increases by 3.6%, and $\Delta$ decreases by 1.7%. These findings suggest that our tests can address the binary classification challenge of detecting a model’s contamination, and more importantly, they can reflect varying degrees of contamination.
5.3 Training Set Leakage Leads to Unfairness
We investigates whether training set leakage leads to evaluation bias. As $D_{\text{tuning}}$ already includes the training set of COCO2017, we remove that portion from $D_{\text{tuning}}$ to obtain LLaVA-1.5-7B-no-coco, thereby simulating a scenario where the model hasn’t encountered the training set.
Model CR PCR $\Delta$
LLaVA-1.5-7B-cont 64.3 63.8 -0.5
LLaVA-1.5-7B-no-cont 61.4 61.5 0.01
Table 4: Detection results on the ScienceQA test set. LLaVA-1.5-7B-cont was trained on 2000 training samples from ScienceQA.
Model CR PCR $\Delta$
LLaVA-1.5-7B-cont 38.1 34.9 -3.2
LLaVA-1.5-7B-no-coco 32.5 31.9 -0.6
Table 5: Detection results on the COCO-Caption2017 validation set. LLaVA-1.5-7B-no-cont: does not include benchmark data but contains the training set from COCO2017.
Tables 5 and 5 demonstrate that training set leakage substantially enhances model performance on test or validation sets, with an average increase of 4.3% in CR and 2.7% in PCR. These findings imply that training using the training set could potentially lead to disproportionate performance advantages.
Takeaways Both training and test set leakage can result in unfairness, and the degree of contamination can be detected through MM-Detect.
@@ 6 At Which Stage is Contamination Introduced?
In this section, we will explore the source of contamination. The training data for most of the MLLMs we examined is openly stated, which prompts the question: If the contamination did not originate from the multimodal training phase, could it have been introduced during the unimodal training phase? In this context, unimodal contamination refers to instances where the text inputs or labels from the benchmark data have already been exposed during the pre-training phase of the LLMs utilized by these models (Section 2.1). To explore this possibility, we examined the LLMs used by the previously tested MLLMs and performed a series of experiments.
@@ 6.1 A Heuristic Experiment for Unimodal Contamination Detection
We employed a heuristic approach based on the intuition that if the LLM can correctly answer an image-required question without the image, it may indicate the leakage of that instance.
@@
The benchmark for this experiment was MMStar, which inherently consists of image-text related questions. The models tested include LLaMA2-7b Touvron et al. (2023b) used by LLaVA-1.5 and VILA, Qwen-7B Bai et al. (2023a) used by Qwen-VL, Mistral-7B-v0.1 Jiang et al. (2023) used by idefics2, Phi-3-small-128k-instruct Abdin et al. (2024) used by Phi-3-vision, Yi-6B AI et al. (2024) used by Yi-VL, and Internlm2-7B Cai et al. (2024) used by InternVL2. To reduce the possibility of the model making random guesses, we appended the prompt, If you do not know the answer, output I don’t know, to the instructions. The metric we reported was the ContRate, representing the frequency these models correctly answered questions without the necessity of image input. It is worth noting that we did not conduct experiments on Fuyu-8b and proprietary models since they did not disclose the unimodal language model and training data used.
@@
Table 6: Contamination rates of the LLMs used by multimodal models. ILM denotes the IL of the corresponding MLLMs.
From Table 6, we can observe that LLaMA2-7b and Qwen-7B still exhibit high contamination rates even without image inputs, with LLaMA2-7b’s rate reaching 25.6%. This indicates that the pre-training corpora of these LLMs may have mixed in text inputs or labels from multimodal benchmarks.
More importantly, we find that most of these LLMs with higher contamination levels correspond to multimodal models like VILA1.5-3B, which has a significant $\Delta$, and LLaVA-1.5-7B and Qwen-VL-Chat, which exhibit high $IL$. This suggests that the contamination suffered by these three MLLMs may not solely originate from the multimodal training phase but rather from the pre-training phase of the LLMs.
6.2 A Review of Visual Instruction Tuning Data for Cross-modal Contamination Detection
@@ 7 Conclusion and Future Work
In this study, we introduce and validate a multimodal data contamination detection framework, MM-Detect, providing new perspectives for evaluating contamination in MLLMs. We discovered that popular multimodal models exhibit varying degrees of data contamination, which directly impacts their performance and generalization ability. In addition, our experimental indicates that MM-Detect is sensitive to varying degrees of contamination and can highlight significant performance improvements due to leakage in the multimodal benchmark training set. Furthermore, we found that the contamination in multimodal models may not solely originate from the cross-modal contamination but could also stem from the unimodal contamination.
@@
We acknowledge several limitations in detecting test set contamination. First, this work is limited to discussions around visual modalities, and does not yet cover other modalities such as audio or video. Second, we only selected widely used and representative multimodal datasets for detection, including multiple-choice datasets and caption datasets, without testing additional datasets. However, we speculate that the method Slot Guessing for Perturbation Caption may also apply to other types of Image Feature Analysis benchmarks.
@@ A.2 Masking-base
These methods involve masking phrases or sentences and providing data from the benchmark to guide the model in filling in the missing parts. However, multimodal datasets often contain images that include the masked portions of sentences, effectively providing answers to the model. This results in significantly higher success rates for multimodal models in predicting missing parts compared to unimodal language models, leading to exaggerated contamination detection. As shown in Table 9, LLaVA-1.5-13b has a high probability of Exact Match for predicting the masked word.
@@ A.3 Comparison-base
These methods identify contamination by comparing the similarity between model outputs and benchmark data. However, MLLMs often undergo data augmentation, causing their outputs to diverge significantly from the labels in benchmark data, making effective contamination detection challenging. From Table 10, we can see that CDD indicates a contamination level of 0% across all multimodal benchmark datasets.
Dataset Contamination Level Split
@@
Appendix B Details of Calculating Atomic Metrics
We formulate the detection pipeline algorithm utilized in MM-Detect, as shown in Algorithm 1.
Algorithm 1 Calculation of Atomic Metrics
1:  Input: A benchmark $D$ with a sequence of datapoints $d_{1},d_{2},\dots,d_{N}$
2:  for $i=1$ to $N$ do
3:     The corresponding detecting method
4:  end for
5:  Compute $\Delta=PCR-CR$, $X$
6:  if $\Delta>-1$ then
7:     Compute $IL=\frac{X}{|D|}$
8:     return $IL$
9:  else
10:     return $\Delta$
11:  end if

### v2

@@ 5 Experiment
Metric CR PCR $\Delta$ $\Phi$ CR PCR $\Delta$ $\Phi$ CR PCR $\Delta$ $\Phi$
@@
Claude-3.5-Sonnet 70.3 65.0 -5.3 15.3 67.3 64.9 -2.4 12.4 36.3 36.4 0.1 15.9
Table 3: Comparison of MLLMs on multi-choice datasets. Bold values represent the most significant $\Delta$ or $\Phi$; color codes denote contamination degree: green for $\mathcal{C}_{\text{Minor}}$, yellow for $\mathcal{C}_{\text{Partial}}$, and red for $\mathcal{C}_{\text{Severe}}$.
In this section, we demonstrate the practicality of our methodology in verifying the leakage of multimodal benchmark datasets across several MLLMs.
5.1 Setup
Models.
We conducted extensive evaluations on nine open-source MLLMs, including LLaVA-1.5-7B Liu et al. (2023a), VILA1.5-3B Lin et al. (2023), Qwen-VL-Chat Bai et al. (2023b), fuyu-8b [^3] https://www.adept.ai/blog/fuyu-8b, idefics2-8b Laurençon et al. (2024), Phi-3-vision-128k-instruct Abdin et al. (2024), Yi-VL-6B AI et al. (2024), InternVL2-8B Chen et al. (2023, 2024b), DeepSeek-VL2-Tiny Wu et al. (2024), as well as three proprietary MLLMs: GPT-4o-2024-08-06 OpenAI (2023), Gemini-1.5-Pro-002 Reid et al. (2024), and Claude-3.5-Sonnet-2024-06-20444https://www.anthropic.com/news/claude-3-5-sonnet.
Benchmark Datasets.
Our analysis leverages two multi-choice datasets: ScienceQA Lu et al. (2022) and MMStar Chen et al. (2024a), along with three caption datasets: COCO-Caption2017 Lin et al. (2015), NoCaps Agrawal et al. (2019), and Vintage [^5] https://huggingface.co/datasets/SilentAntagonist/vintage-artworks-60k-captioned. MMStar and Vintage, owing to their recent inception, serve to contrast leakage levels with other datasets. We randomly selected 2000 and 1340 samples from ScienceQA’s training and test sets, respectively, with 1000 samples from the other datasets. Given the unavailability of public test labels for COCO-Caption2017 and NoCaps, we used their validation sets.
5.2 Main Results
@@
Metric CR PCR $\Delta$ $\Phi$ CR PCR $\Delta$ $\Phi$ CR PCR $\Delta$ $\Phi$
@@
idefics2-8b 43.5 42.3 -1.2 21.2 42.6 37.5 -5.1 23.3 18.5 17.0 -1.5 14.5
Phi-3-vision-128k-instruct 38.8 39.3 0.5 19.4 36.9 33.3 -3.6 19.7 17.4 11.7 -5.7 14.3
Yi-VL-6B 43.9 43.3 -0.6 19.4 37.2 36.1 -1.1 17.5 3.3 4.2 0.9 2.8
InternVL2-8B 53.3 51.9 -1.4 20.4 48.0 46.2 -1.8 20.9 28.0 28.7 0.7 18.8
DeepSeek-VL2-Tiny 23.8 21.4 -2.4 13.5 19.3 18.1 -1.2 12.2 7.5 6.9 -0.6 6.3
@@
GPT-4o 58.1 54.4 -3.7 23.1 54.2 55.1 0.9 19.4 36.3 38.4 2.1 20.1
Gemini-1.5-Pro 57.5 55.3 -2.2 21.6 51.2 52.0 0.8 18.7 46.3 41.0 -5.3 28.3
Claude-3.5-Sonnet 53.7 51.0 -2.7 21.8 50.8 51.5 0.7 20.0 35.2 33.0 -2.2 21.3
Table 4: Comparison of MLLMs on caption datasets.
@@
Table 3 yields several conclusions: (1) Both open-source and proprietary models exhibit contamination. For example, on the ScienceQA training set, both open-source models like LLaVA-1.5-7B and idefics2-8b and proprietary model Gemini-1.5-Pro show minor contamination degree. (2) Proprietary models are more contaminated. Claude-3.5-Sonnet, for instance, registers a severe $\Delta$ with higher $\Phi$ values on both ScienceQA training and test sets, indicating extensive leakage. (3) Training set leakage is more pronounced than test set leakage. On the ScienceQA dataset, models generally exhibit larger $\Delta$ values in the training set, for instance, Claude-3.5-Sonnet shows $\Delta=-5.3$ on training versus $\Delta=-2.4$ on the test set, while most models have near-zero $\Delta$ on the test set. (4) Older benchmarks are more prone to leak. The older ScienceQA test set shows more severe leakage (e.g., Claude-3.5-Sonnet’s $\Delta$ of -2.4) compared to the newer MMStar validation set (e.g., fuyu-8b’s $\Delta$ of -1.2).
@@
Table 4 yields several conclusions: (1) Both open-source and proprietary models exhibit contamination on caption datasets. For example, in the COCO Validation Set, open-source models such as DeepSeek-VL2-Tiny and proprietary models like GPT-4o record a significant contamination degree. (2) Leakage levels vary significantly by benchmark. For example, on the NoCaps Validation Set, open-source models exhibit more pronounced contamination degree than proprietary models, whereas the trend reverses on the COCO Validation Set. These findings confirm that caption datasets are vulnerable to leakage, with proprietary models generally exhibiting more pronounced contamination effects.
@@ 6 At Which Stage is Contamination Introduced?
In this section, we investigate the source of contamination in MLLMs. Although the training data for some MLLMs is openly documented, an important question remains: if contamination does not arise during the multimodal training phase, could it stem from the unimodal (pre-training) phase, as defined in §2.1? To address this possibility, we examined the underlying LLMs of the evaluated MLLMs and conducted a series of experiments (§6.1). We also explored the origins of cross-modal contamination arising during visual instruction tuning (§6.2).
@@ 6.1 A Heuristic Experiment for Unimodal Contamination Detection
We employed a heuristic approach based on the intuition that if an LLM can correctly answer an image-required question without the image, it may indicate the leakage of that instance.
@@
We used MMStar as the benchmark, where every question relies on visual input for correct answers. The tested models include LLaMA2-7B Touvron et al. (2023b) (used by LLaVA-1.5 and VILA), Qwen-7B Bai et al. (2023a) (used by Qwen-VL), Mistral-7B-v0.1 Jiang et al. (2023) (used by idefics2), Phi-3-small-128k-instruct Abdin et al. (2024) (used by Phi-3-vision), Yi-6B AI et al. (2024) (used by Yi-VL), and Internlm2-7B Cai et al. (2024) (used by InternVL2). To discourage random guessing, we appended the prompt “If you do not know the answer, output I don’t know” to the instructions. Accuracy, the frequency with which models correctly answer questions without image input, is reported as the primary metric. Note that we did not evaluate Fuyu-8B and proprietary models since their unimodal LLM and training data remain undisclosed.
@@
Table 5: Contamination rates of LLMs used by MLLMs. $\Phi$M denotes the $\Phi$ of the respective MLLMs.
Table 5 yields several conclusions: (1) Contamination occurs in LLM. All models exhibit varied contamination rates, indicating that their pre-training data likely included text from multimodal benchmarks. (2) Elevated LLM contamination correlates with increased MLLM leakage. For instance, VILA1.5-3B and Qwen-VL-Chat exhibit significant $\Phi$ values that mirror their underlying LLM contamination levels. These findings suggest that contamination in these MLLMs may originate partly from the LLMs’ pre-training phase, rather than solely from multimodal training.
6.2 Tracing Origins: A Review of MLLM’s Visual Instruction Tuning Data
@@ 7 Conclusion and Future Work
In this study, we introduce and validate a multimodal data contamination detection framework, MM-Detect, providing new perspectives for evaluating contamination in MLLMs. We discovered that popular MLLMs exhibit varying degrees of data contamination, which directly impacts their performance and generalization ability. In addition, our experiment indicates that MM-Detect is sensitive to varying degrees of contamination and can highlight significant performance improvements due to the leakage in the multimodal benchmark training set. Furthermore, we found that the contamination in MLLMs may not solely originate from the cross-modal contamination but could also stem from the unimodal contamination.
@@
We acknowledge several limitations in our work. First, this work is limited to discussions around visual modalities, and does not yet cover other modalities such as audio or video. Second, we only selected widely used and representative multimodal datasets for detection, including multiple-choice datasets and caption datasets, without testing additional datasets, such as open-ended generation and cloze questions. However, we speculate that the method Slot Guessing for Perturbation Caption may also apply to other types of image-feature-analyzing benchmarks. Third, the effectiveness of Option Order Sensitivity Test can be undermined by option shuffling, which, while potentially improving model performance, is computationally expensive and may increase the training cost.
@@ A.2 Masking-base
These methods involve masking phrases or sentences and providing data from the benchmark to guide the model in filling in the missing parts. However, multimodal datasets often contain images that include the masked portions of sentences, effectively providing answers to the model. This results in significantly higher success rates for MLLMs in predicting missing parts compared to unimodal language models, leading to exaggerated contamination detection. As shown in Table 8, LLaVA-1.5-13b has a high probability of Exact Match for predicting the masked word.
@@ A.3 Comparison-base
These methods identify contamination by comparing the similarity between models’ outputs and benchmark data. However, MLLMs often undergo data augmentation, causing their outputs to diverge significantly from the labels in benchmark data, making effective contamination detection challenging. From Table 9, we can see that CDD (Contamination Detection via Output Distribution) indicates a contamination metric of 0% across all multimodal benchmark datasets.
Dataset Contamination Metric Split